Accueil > Sites heberges > CFHTLS-Deep > 2003 runs

 

2003 runs

Par Genevieve Soucail - 24/04/2006

 

Summary of 2003 runs

  • Summary of queue run 03BQ06 : Dec. 16 - Dec. 31
    • From Dec. 15 to 26, rather good observations, with photometric sky most of the time, but very variable and unstable seeing, ranging from 0.6" to more than 1.5" !!! A significant fraction of the observations of the Deep was not validated, although executed for the Supernovae programme. Then the last 4 nights were lost because of bad weather.
    • CCD#03 not working.
    • D1 observed in r’i’z’ only for a total of 5.3 hours.
    • Observations of D2 started in 5 filters for a total of 3.5 hours.

  • Observations of D4 are closed (December 16, 2003). See the D4 web page for details.

  • Summary of queue run 03BQ05 : Nov. 17 - Dec. 03
    • A terrible run with more than 65 % of the run lost to weather. An even higher fraction for the Deep survey.
    • 1h30 of validated exposures for D4 and 2h37 for D1. No observations started on D2.

  • Summary of queue run 03BQ04 : Oct. 16 - Nov. 01
    • A difficult run because of weather variability and periods of thick clouds and strong absorption. Again CCD#03 stopped working by Oct. 27 until the end of the run, affecting only a few exposures for the Deep.
    • Variable seeing, with exposures not validated on Oct. 16 and 17. It then improved significantly for the acquired images.
    • Grosso modo, 3 epochs of observation in griz for D1 and D4, plus a few additional exposures. And one serie of u-band images for D1 and 2 series for D4, compensating the previous run when no u-data was obtained.
    • In total, D4 observed for 9.1 hours and D1 for 8.6 hours.

  • The coordinates of D2 have changed !!! The CFHTLS/D2 will move to the COSMOS/ACS survey area (field observed by one of the treasury programmes accepted on HST). The main advantage is the rapid availability of a large set of multi-wavelength follow-ups, from the X-rays (XMM) to the UV (GALEX) and the far-IR (SIRTF), as well as multi-objects spectroscopy with VIMOS at ESO.

  • Summary of queue run 03BQ03 : Sep 18 - Oct 4
    • A pretty good run for the first half of the run, then strongly limited by some serious technical problems with the instrument : due to a serious problem with the filter jukebox, it was dismounted on Sep. 25. Only one filter available per night during
    • CCD#03 was not working all along the run.
    • Very good seeing all along the run. 
    • D4 observed for a total of 8.4 hours . Mars was very close to the field (less than one degree) and created strange features on some parts of the mosaic. About 3 CCDs losts (#10-19-28) on Sep. 21 and (#12-21-30) on Sep. 25.
    • D1 observed for a total of 9.1 hours.

  • Summary of queue run 03BQ02 : Aug20 - Sep4 :
    • Not a bad run indeed for the Deep survey !
    • D4 (RA=22h) was intensively observed (not far from nominal observations, indeed !) with a total integration time of 10.4 hours.
    • D1 was observed for a few hours only (3.5 hours), as it was observable only 1 to 2 hours per night (RA=2h). Only i and r observations (+ some z ones) were obtained, necessary to build a reference field for the SN searches. More is expected for next run.

  • Summary of 2003A : Difficult semester, with significant difficulties all along the runs :
    • Before the April run (03AQ04), we can hardly consider that the image quality was sufficient enough. However, the data can be used for photometric calibration, and was used in order to build the reference field for the SN search 
    • From May 21 to 29 (included), only half of the mosaic was used (CCDs 00 to 17), because of the failure of the second CCD controller. Each pointing was duplicated with a shift corresponding to half the mosaic. Be careful when the images will be stacked.
    • The July run was nearly completely washed out by poor weather ! 
    • In total, more than 11 hours of integration was validated on the field, i.e. within specs of the Deep requirements (seeing < 0.9", acceptable up to 1", transparency better than 90%)
    • However the distribution among the 5 filters is not quite optimal. No u* data were acquired, z’ data are useless and not validated (lack of fringe frames to correct the fringe pattern) and "too much" i’ data is available compared to the other bands. Look at the relative distribution among the 5 filters

  • The adjustment of the focal plane of Megaprime has been fixed for the semester 2003B (July 23, 2003).
 
Back to recent updates

 

 

[Page Précédente] [Dans la même rubrique] [Sommaire]

 

 

Site du CNRS Logo OMP
Logo UPS
Logo INSU-CNRS

Annuaire

Rechercher

Sur ce site



Bibliotheque

Laboratoire Astrophysique de Toulouse - Tarbes (UMR5572)

CNRS (Midi-Pyrénées)

Univ. Paul Sabatier

Liens utiles

Accueil Imprimer Contact mail Plan du site Crédits